OpenStudy (bgrg007):

CALC 3 Q Help!! Reduce the equation to one of the standard forms and classify the surface. x^2-y^2+z^2-2x+2y+4z+2=0

OpenStudy (anonymous):

complete the square for all variables. \[x^2 - 2x = (x-1)^2 - 1\]\[y^2 + 2y = (y+1)^2 - 1\]\[z^2 + 4z = (z+2)^2 - 4\] so we have\[(z+2)^2 = 4 - (x-1)^2 - (y+1)^2\] solve for z = f(x,y) let me know if i should keep going

OpenStudy (anonymous):

my bad it's 2y - y^2 = (1 - y)^2 - 1

OpenStudy (anonymous):

(x-1)^2-(y-1)^2+(z+2)^2=3 

OpenStudy (anonymous):

i havent drawn surfaces in a while so it's probably best the way vijay has it, which is, not solving for z = f(x,y). that would be a useless step in the drawing process and probably identifying process

OpenStudy (bgrg007):

i got (z+2)^2=4-(x-1)^2+(y-1)^2

OpenStudy (anonymous):

that is the right answer

OpenStudy (bgrg007):

so im not going crazy. haha the answer i found on chegg was completely different so i was worried. also, the drawing @vvijay has would be a hyperboloid with 2 sheets?